# On the attribution of weather events to climate change using a fit to extreme value distributions

### v2

@@ 3.4 Madagascar Drought
Finally, we consider the attribution analysis for the 2019–2021 drought in Madagascar, following the lead of HW2022. This case differs from the two examined above in that it analyzes precipitation extremes that are positive by definition. HW2022, therefore, used a different extreme value distribution function, the GPD, which is meant to represent the high tail of the distribution of extreme events over a specified threshold Coles et al. (2001). As discussed in Section 2, HW2022 analyzed the driest 20% of the 2-year averages of precipitation data from 1951–2020 multiplied by minus one to turn minima (droughts) into maxima. We follow their example, with the important caveat made in Section 22.2 that the validity of using a GPD this way for precipitation data is not guaranteed given that there is a hard maximum of zero for these data of negative precipitation rates, a hard maximum that seems inconsistent with the definition of the GPD. We also note that taking the lowest 20% of 2-year precipitation data points amounts to only 14 data points, likely too low a number for a reliable attribution.
@@
Fig. 8d shows the MRL plot (section 22.5) for the Madagascar Drought case. As a reminder, this plot needs to scale linearly within the range of the chosen threshold value for the GPD fit (section 22.5) in order for the GPD analysis to be justified and self-consistent. For small and large values of precipitation, the MRL curve appears to curve nonlinearly, indicating that a GPD fit is likely invalid using those thresholds. For precipitation thresholds ranging from roughly 1.6 to 2 mm/day, the MRL curve plateaus. The orange line shows the expected MRL line, ($\sigma+u\xi/(1-\xi)$); that slope does not align with the MRL curve around the value of 1.9 corresponding to the 20th percentile precipitation threshold chosen by Harrington et al. (2022). Likely the small number of data points is an obstacle for a reliable MRL plot here. In general, we suggest that this MRL analysis should be helpful in examining the self-consistency of applying a GPD to precipitation data. In this particular case it is not obvious if the required linear range exists in this plot to a degree that justifies using a GPD distribution and allows selecting a threshold.
@@
We find again that an extreme value distribution is not required by the data distribution, and a GMST dependence is not needed: the GPD fit to the bottom 20th percentile precipitation data is no better than a log-normal distribution fitted to all of the precipitation data, both with constant distribution parameters that do not depend on the GMST (Fig. 8a). A log-normal distribution seems to be a reasonable null hypothesis for the positive precipitation data. In fact, the empirical CDF is practically unconstrained at its lower tail because there are no data points below a value of 1.25 mm/day.
@@
Figure 8: Madagascar Drought: (a) Evaluating the need for a GPD fit. The empirical CDF is shown (red line, with 90th and 95th percentile confidence intervals shown by red shadings), based on 24-month running means of daily averaged annual precipitation over the Madagascar area defined in HW2022 using ERA5 precipitation. A GPD fit to the bottom 20th percentile precipitation (black) and a log-normal fit to the full distribution (blue). (b) The deviance statistic between a GPD with and without a GMST component in the CESM Large Ensemble (blue) and observations (green). (c) Scatter plot of the 2-year running mean of ERA5 daily precipitation over the Madagascar area defined in HW2022 against the 4-year running mean of annual GMST for the lowest 20 percent precipitation data. The dashed blue line shows the exponential GPD log-likelihood fit for the threshold parameter for the observations without resampling (eqn. 7) following HW2022. (d) MRL plot (section 22.5) for different precipitation thresholds. The blue shading defines a region of confidence calculated as $\pm$ one standard deviation. The red line denotes the threshold used in the GPD fit of HW2022 (i.e., the 20th percentile). The orange line represents the full linear equation from Equation 9 ($\sigma+u\xi/(1-\xi)$).
@@
The deviance statistic for the observations is marginally above the red significance line (leftmost green bar, Fig. 8b) suggesting that the inclusion of the additional model parameter $\alpha$ may be justified (uncertainty plots for the GPD parameters for observations are indicated in Supplementary Fig. SI-6). However, the bootstrap estimates for this parameter shown in Supplementary Fig. SI-7a show it to range over both negative and positive values, indicating that it cannot be calculated with any confidence. None of the 35 ensemble members exceed the deviance statistic significance level. It could be argued that Figure 1c, showing the analysis of the pre-industrial run, is evidence of the absence of any strong effect of natural variability modes on both GMST and Madagascar drought, in contrast with the analysis of Australian temperatures in Figure 1a. Also, Figure 8b may suggest that the CESM Large Ensemble does not adequately capture the effect of anthropogenic climate change on droughts in Madagascar.
@@
As a reminder, the threshold parameter of the GPD, corresponding to the maximum precipitation value considered as an extreme drought year in this attribution case, was assumed by HW2022 to vary exponentially with the GMST ($\overline{T}$), as $\mu=\mu_{0}\exp(\alpha\overline{T}/\mu_{0})$ with $\mu_{0}<0$. If droughts become more severe with increasing GMST, we expect the threshold parameter for the distribution of the negative precipitation to get larger (less negative) with increasing GMST, and therefore $\alpha/\mu_{0}$ to be negative and thus $\alpha$ to be positive. The value of ${\alpha}_{\mathrm{GPD}}$ calculated from observations is small and indistinguishable from zero (Fig. 8c). The exponential fit was assumed by HW2022 in order to ensure that the scaling factor $\exp(\alpha\overline{T}/\mu_{0})$ applied to both $\mu_{0}$ and $\sigma_{0}$ is strictly positive, so the distribution has fixed dispersion Philip et al. (2020). A scatter plot of ERA5 precipitation against GMST anomalies is shown in Fig. 8c. The dashed blue line indicated in the plot represents the exponential fit using the $\alpha$ and $\mu_{0}$ coefficients from the maximum log-likelihood GPD fit. This fit looks notably linear, consistent with the Taylor expansion of the exponential function,
$\mu=\mu_{0}\exp(\alpha\overline{T}/\mu_{0})\approx\mu_{0}(1+\alpha\overline{T}/\mu_{0})=\mu_{0}+\alpha\overline{T}.$
@@
The small value of $\alpha$ found for observations indicates that the increase in GMST has had little effect on the statistics of two-year-average precipitation in Madagascar. Whereas this is also the conclusion reached by HW2022, we found that the data can be fitted with a log-normal distribution rather than GPD and without GMST dependence (Fig. 8a). Further, the deviance statistic only marginally justifies adding GMST dependence for observations and does not justify adding such dependence for model output (Fig. 8b). The fact that only 14 data points are involved in the fit seems much too small to draw any meaningful conclusions on an extreme value distribution as it leads to a large uncertainty interval shown by the shading. Supplemental Fig. SI-8a indicates that even if using the equivalent of 35 times as much data, the uncertainty range for $\alpha$ includes both positive and negative values. One could, therefore, conclude that the GPD approach is not likely to lead to confident attribution in this case in the next decades and beyond—even if there were actually an anthropogenic signal—given the small sample sizes from sub-sampling the bottom 20 percent of the precipitation data.
@@ 4 Conclusions
We were inspired in this work by a series of attribution studies by the “World Weather Attribution Project” (WWA) and evaluated the part of their methodology that involves a fit of observations to extreme value distribution functions. The full protocol used by this effort Philip et al. (2020) uses additional tools and model analysis. The WWA method involves fitting an extreme value distribution function to an observed record, with the distribution parameters depending on global mean surface temperature (GMST), whose variation is known to represent anthropogenic climate change. For example, the location parameter of an extreme value distribution (GEV) is made a linear function of GMST, $\overline{T}$, as $\mu=\mu_{0}+\alpha\times\overline{T}$. The GMST dependence $\alpha$ is then used to calculate the change in return time of extreme events. This method was applied to many extreme events such as heatwaves, cold spells, floods, droughts, and more (e.g., van Oldenborgh et al., 2016; van der Wiel et al., 2017; van Oldenborgh et al., 2017; Ciavarella et al., 2021; van Oldenborgh et al., 2021; Harrington et al., 2022). We note that we are not the first study to scrutinize the WWA approach. García-Portela and Maraun (2023) discussed the WWA approach for CMIP model selection, i.e., that model spread should appropriately represent the full spread of dynamical aspects rather than subsetting to specific climate models that perform well on present-day climate. In our case, we examined the robustness of the results of three example WWA studies, demonstrated when they can fail due to the effects of natural variability, and suggested several additional statistical tests, reviewed below, that may increase confidence in the results of such an analysis.
@@
We began by examining the GMST dependence of extreme heatwave events in Siberia (following CC2021), Australia (following OK2021), and drought in Madagascar (following OK2021), using data from a CESM preindustrial simulation. We found a significant GMST dependence in the first two cases, indicating that such dependence cannot be used to unequivocally identify the effects of anthropogenic climate change. The GMST dependence, in these cases, is a result of regional natural variability such as ENSO, which affects the extreme events being examined, and also weakly affects GMST. This seems to invalidate the use of a non-zero $\alpha$ to quantify the effect of anthropogenic climate change. The value of $\alpha$ in the Australian case is especially large, seemingly indicating that every degree increase in GMST leads to 5-degree increase in Australian heatwaves. Instead, we inferred that the large values of $\alpha$ are a result of natural variability (likely ENSO) affecting GMST by 0.2 degrees for every 1-degree effect it has on Australian heatwaves.
@@
We tested if an extreme value distribution is indeed needed by plotting the empirical cumulative distribution function (CDF) with bootstrapping error bars (Figs. 2a, 6a, 8a). We found that fits using a standard distribution (normal or log-normal) fall within the error bars of the empirical distribution, indicating that there may not be a need for an extreme value distribution. Similarly, the standard distributions seem to fit the data well without a GMST dependence, indicating that such dependence may not be justified by the data. We further tested if the addition of a GMST dependence of the distribution parameters is justified statistically using the log-likelihood ratio test, also known as the deviance statistics. When adding a parameter (specifically $\alpha$), one expects the fit to improve, but the question is whether the improvement is beyond what is expected simply due to the addition of one more model parameter, so that one is convinced that the additional parameter represents a genuine effect in the data. We found that in the Australia and Madagascar cases, there is no justification for the addition of $\alpha$.
@@
The Madagascar drought attribution (HW2022) was based on only fourteen of the lowest two-year-averaged precipitation rates, hardly sufficient to construct a meaningful empirical CDF or to constrain the four parameters of the GPD distribution used for that purpose. In addition, computing two-year averages of precipitation and applying an extreme value distribution may lead to inconsistencies a priori given that averages converge to a Gaussian distribution (under the Central Limit Theorem). The GPD fit to the record, in this case, is meant for high values above a threshold, and the expected mean needs to vary linearly with the threshold chosen. We suggested that a mean residual life (MRL) plot can generally be used to test the self-consistency of using a GPD. In this case, it was not clear that these drought events are well represented by a GPD, but certainly, the number of data is insufficient for a meaningful MRL analysis in this particular case. To deal with the loss of data associated with this approach and difficulties in choice of the optimal precipitation threshold, Naveau et al. (2016) suggest instead using a statistical model fit to the full distribution of precipitation data, which has the added benefit of following the rules of extreme value theory for both high and low precipitation events. While our results—that attribution of two-year-averaged precipitation is not possible—agree with those of HW2022, we emphasized our different perspective that it is unlikely that attribution would be possible based on such a small number of data points, as well as the other difficulties pointed out above. In contrast, other approaches evaluating changes in the seasonal cycle of precipitation have led to the attribution of shifts in seasonal rainfall patterns in Southern Madagascar to climate change Rigden et al. (2024).
@@
An important issue raised by this work is the discrepancy between the modeled and observed findings, as seen, for example, in the deviance statistic results for the observations versus CESM Large Ensemble (Figs. 2b, 6b, 8b). This may be due to a particularly favorable sequence of internal variability in the observations, measurement error in the observational record, or issues with the simulation of extreme weather in the CESM Large Ensemble. It would be helpful to test if these discrepancies also occur with other model large ensembles, including those used by WWA (Philip et al. (2020), OK2021).
@@
Overall, our results suggest uncertainty in the interpretation of the dependence of the parameters of extreme event distributions on the global mean surface temperature as reflecting anthropogenic climate change. The proposed and demonstrated deviance statistic, MRL plot, and a careful analysis of uncertainty ranges, especially for $\alpha$, may help avoid misguided confidence in attribution results. As noted in the introduction, WWA does not exclusively rely upon the empirical methodology evaluated here in order to reach conclusions, and also evaluates other historical data, model simulations, and the dynamical context. WWA also highlights some caveats associated with their study, including the necessity of testing for other probability laws (e.g., Gaussian, Gamma, etc.), covariates (ENSO, PDO, etc.), among others Philip et al. (2020). Nevertheless, the WWA empirical methodology examined here still figures prominently in their analyses, and we have highlighted several caveats with its interpretation, foremost that internal variability can give rise to the appearance of strong covariance between GMST and local extremes. Our results do not provide evidence against a relationship between GMST and local extremes but, rather, that internal variability needs to be taken into account when fitting a distribution function to the record. As currently implemented, the WWA approach can lead to first-order biases in some cases. It, therefore, seems that it may be best to use the WWA empirical methodology in conjunction with other attribution approaches, which are summarized in the Introduction (e.g., Stott et al., 2016; Knutson et al., 2017; Seneviratne et al., 2021).
@@
We suggest two key takeaways. First, controlling for internal variability, possibly by making the distribution parameters depend on additional covariates such as ENSO, appears important. This improvement upon the current method should be further developed. Second, empirical attribution techniques should be examined by applying the approach to simulations with and without changes to anthropogenic emissions. Application to such simulations will indicate whether the test is capable of discerning anthropogenic influences upon the statistics of extreme events. Although our focus is on the analysis of observation, some of our results are based on model output, and we use only one global climate model (CESM). This analysis should be extended to other climate models as well for purposes of further validation, although it seems that the main point of section 3.1 regarding the interaction of natural variability and global mean surface temperature should be robust to model choice, at least qualitatively.

### v3

@@ 3.4 Madagascar Drought
Finally, we consider the attribution analysis for the 2019–2021 drought in Madagascar, following the lead of HW2022. This case differs from the two examined above in that it analyzes precipitation extremes that are positive by definition. HW2022, therefore, used a different extreme value distribution function, the GPD, which is meant to represent the high tail of the distribution of extreme events over a specified threshold Coles et al. (2001). As discussed in Section 2, HW2022 analyzed the driest 20% of the 2-year averages of precipitation data from 1951–2020 multiplied by minus one to turn minima (droughts) into maxima. We follow their example, with the important caveat made in Section 22.2 that the validity of using a GPD this way for precipitation data is not guaranteed, given that there is a hard maximum at zero precipitation that seems inconsistent with the definition of the GPD. We also note that taking the lowest 20% of 2-year precipitation data points amounts to only 14 data points, likely too low a number for a reliable attribution.
@@
Fig. 8d shows the MRL plot (section 22.5) for the Madagascar Drought case. As a reminder, this plot needs to scale linearly within the range of the chosen threshold value (section 22.5) in order for the GPD analysis to be justified and self-consistent. For small and large values of precipitation, the MRL curve appears to curve nonlinearly, indicating that a GPD fit is likely invalid using those thresholds. For precipitation thresholds ranging from roughly 1.6 to 2 mm/day, the MRL curve plateaus. The orange line shows the expected MRL line, ($\sigma+u\xi/(1-\xi)$); that slope does not align with the MRL curve around the value of 1.9 corresponding to the 20th percentile precipitation threshold chosen by Harrington et al. (2022). In general, we suggest that this MRL analysis should be helpful in examining the self-consistency of applying a GPD to precipitation data, though the small number of data points appears an obstacle for a reliable MRL analysis in this particular case. Specifically, it is not obvious if the required linear range exists to a degree that justifies using a GPD distribution and allows for selecting a threshold.
@@
We find again that an extreme value distribution is not required by the data distribution. The GPD fit to the bottom 20th percentile precipitation data is no better than a log-normal distribution fitted to all of the precipitation data (Fig. 8a). Furthermore, both the log-normal and GPD fit indicate no need of a GMST dependence, in keeping with the small value of $\alpha$ found by HW2022. A log-normal distribution seems a reasonable null hypothesis for the positive precipitation data. In fact, the empirical CDF is practically unconstrained at its lower tail because there are no data points below a value of 1.25 mm/day. Supplementary Fig. SI-6 for GPD, log-normal, and empirical distributions also suggests the log-normal distribution looks reasonable, albeit acknowledging the highly limited sample size in this case.
@@
Figure 8: Madagascar Drought: (a) Evaluating GPD and log-normal fits to rainfall data. An empirical CDF of rainfall is shown (red line, with 90th and 95th percentile confidence intervals shown by red shadings), based on 24-month running means of daily precipitation over the Madagascar area defined in HW2022 using ERA5 precipitation. A GPD fit to the bottom 20th percentile precipitation (black) and a log-normal fit to the full distribution (blue). (b) The deviance statistic between a GPD with and without a GMST component in the CESM Large Ensemble (blue) and observations (green). (c) Scatter plot of the 2-year running mean of ERA5 daily precipitation over the Madagascar area defined in HW2022 against the 4-year running mean of annual GMST for the lowest 20 percent precipitation data. The dashed blue line shows the exponential GPD log-likelihood fit for the threshold parameter for the observations without resampling (eqn. 7) following HW2022. (d) MRL plot (section 22.5) for different precipitation thresholds. The blue shading defines a region of confidence calculated as $\pm$ one standard deviation. The red line denotes the threshold used in the GPD fit of HW2022 (i.e., the 20th percentile). The orange line represents the full linear equation from Equation 9 ($\sigma+u\xi/(1-\xi)$).
@@
The deviance statistic for the Madagascar precipitation observations is marginally significant at the Xlevel (leftmost green bar is above the red significance line , Fig. 8b), indicating that the inclusion of the additional model parameter $\alpha$ is justified. Uncertainty plots for the GPD parameters fit to these precipitation observations are indicated in Supplementary Fig. SI-7. The bootstrap estimates for the $\alpha$ parameter (Supplementary Fig. SI-8a) ranges over both negative and positive values, indicating that it does not significantly differ from the null. None of the 35 ensemble members exceed the deviance statistic significance level. Figure 1c shows that there is no relation between these extreme drought events and GMST. One interpretation is that this is due to the absence of any strong effect of internal variability modes on both GMST and Madagascar drought, in contrast with the analysis of Australian temperatures in Figure 1a. Also, Figure 8b may suggest that the CESM Large Ensemble does not adequately capture the effect of ACC on droughts in Madagascar.
@@
As a reminder, the threshold parameter of the GPD, corresponding to the maximum precipitation value considered as an extreme drought year in this attribution case, was assumed by HW2022 to vary exponentially with the GMST ($\overline{T}$), as $u=u_{0}\exp(\alpha\overline{T}/u_{0})$ with $u_{0}<0$. If droughts become more severe with increasing GMST, we expect the threshold parameter for the distribution of the negative precipitation to get larger (less negative) with increasing GMST, and therefore $\alpha/u_{0}$ to be negative and $\alpha$ to be positive. The value of ${\alpha}_{\mathrm{GPD}}$ calculated from observations is small and indistinguishable from zero (Fig. 8c). The exponential fit was assumed by HW2022 in order to ensure that the scaling factor $\exp(\alpha\overline{T}/u_{0})$ applied to both $u_{0}$ and $\sigma_{0}$ is strictly positive, so the distribution has fixed dispersion Philip et al. (2020). A scatter plot of ERA5 precipitation against GMST anomalies looks notably linear (see the dashed blue line in Fig. 8c), consistent with the Taylor expansion of the exponential function,
$u=u_{0}\exp(\alpha\overline{T}/u_{0})\approx u_{0}(1+\alpha\overline{T}/u_{0})=u_{0}+\alpha\overline{T}.$
@@
The small value of $\alpha$ found for observations indicates that the increase in GMST has had little effect on the statistics of two-year-average precipitation in Madagascar. Whereas this is also the conclusion reached by HW2022, we found that the data can be fitted with a log-normal distribution rather than GPD and without GMST dependence (Fig. 8a). Further, the deviance statistic only marginally justifies adding GMST dependence for observations and does not justify adding such dependence for model output (Fig. 8b). The fact that only 14 data points are involved in the fit seems much too small to draw any meaningful conclusions on an extreme value distribution. Supplementary Fig. SI-9a indicates that even if using the equivalent of 35 times as much data, the uncertainty range for $\alpha$ includes both positive and negative values. One could, therefore, conclude that there is no clear relationship between GMST and low rainfall in this region.
@@ 4 Conclusions
We were inspired in this work by a series of attribution studies by the “World Weather Attribution Project” (WWA) and evaluated the part of their methodology that involves a fit of extreme value distribution functions to observations. The full protocol used by this effort Philip et al. (2020) involves additional tools and model analyses, including selecting and defining extreme events, analyzing trends in observational data, using multi-model ensembles to simulate event scenarios with and without human-induced climate change, performing quality checks on model accuracy, and synthesizing results to determine climate change’s influence on the event’s intensity and likelihood. A central WWA method nevertheless involves fitting an extreme value distribution function to an observed record, with the distribution parameters depending on global mean surface temperature (GMST), whose variation is taken to represent ACC. For example, the location parameter of an extreme value distribution (GEV) is sometimes made a linear function of GMST, $\overline{T}$, as $\mu=\mu_{0}+\alpha\times\overline{T}$. The GMST dependence included in the function is then used to calculate the change in return time of extreme events. This method was applied to many extreme events such as heatwaves, cold spells, floods, droughts, and more (e.g., van Oldenborgh et al., 2016; van der Wiel et al., 2017; van Oldenborgh et al., 2017; Ciavarella et al., 2021; van Oldenborgh et al., 2021; Harrington et al., 2022). We examined the robustness of the results of three example WWA studies, demonstrated when the results may be unreliable due to the effects of internal variability, and suggested several additional statistical tests, reviewed below, that may increase confidence in the results of such an analysis.
@@
We began by examining the GMST dependence of extreme heatwave events in Siberia (following CC2021) and Australia (following OK2021), as well as drought in Madagascar (following OK2021), using data from a CESM preindustrial simulation. We found a significant GMST dependence in the first two cases, indicating that such dependence can result from internal variability rather than only represent the effects of ACC. The GMST dependence, in these cases, is a result of regional internal variability such as ENSO or other slowly varying modes, which affects the extreme events being examined, and also weakly affects GMST. This points that the value of $\alpha$ calculated in the more realistic case where both effects are present includes the effects of internal variability in addition to those of ACC. The value of $\alpha$ in the Australian case is especially large, and could naively be interpreted as indicating that every degree increase in GMST is associated with a 5-degree increase in Australian heatwaves. Heuristically, the large values of $\alpha$ can be thought of as a result of internal variability (possibly ENSO) affecting GMST by 0.2 degrees for every 1-degree effect it has on Australian heatwaves. As noted earlier, this issue could be mitigated in part by including indices for modes of internal variability as covariates in the model fitting.
@@
We tested if an extreme value distribution is indeed needed by plotting the empirical cumulative distribution function (CDF) with bootstrapping error bars (Figs. 2a, 6a, 8a). We found that fits using a standard distribution (normal or log-normal) fall within the error bars of the empirical distribution, indicating that the observations do not empirically allow for distinguishing between the characteristics of extreme value and normal distributions. The differing tails of these distributions, however, lead to different extreme event return times. Similarly, the standard distributions seem to fit the data well without a GMST dependence, indicating that such dependence may not be justified by the data. We further tested if the addition of a GMST dependence of the distribution parameters is justified statistically using the log-likelihood ratio test, also known as the deviance statistics. When adding a parameter — in this case $\alpha$ — one expects the fit to improve, but the question is whether the improvement is beyond what is expected simply due to the addition of one more model parameter, so that one is convinced that the additional parameter represents a genuine effect in the data. We found that in the Australia and Madagascar cases, there is no justification for the addition of $\alpha$.
@@
The Madagascar drought attribution (HW2022) was based on only fourteen of the lowest two-year-averaged precipitation rates, hardly sufficient to construct a meaningful empirical CDF or to constrain the four parameters of the GPD distribution used for that purpose. In addition, computing two-year averages of precipitation and applying an extreme value distribution may lead to inconsistencies a priori given that averages converge to a Gaussian distribution (under the Central Limit Theorem). The GPD fit to the record, in this case, is meant for high values above a threshold, and the expected mean needs to vary linearly with the threshold chosen. We follow Coles et al. (2001) in suggesting that a mean residual life (MRL) plot can generally be used to test the self-consistency of using a GPD. In this case, it was not clear that these drought events are well represented by a GPD, but with the aforementioned overarching concern regarding the number of data points. To deal with the loss of data associated with this approach and difficulties in choice of an optimal precipitation threshold, Naveau et al. (2016) suggest instead using a statistical model fit to the full distribution of precipitation data, which has the added benefit of following the rules of extreme value theory for both high and low precipitation events. While our results—that attribution of two-year-averaged precipitation is not possible—agree with those of HW2022, we emphasized our different perspective: it is unlikely that attribution would be possible based on such a small number of data points, as well as the other difficulties pointed out above. In contrast, other approaches evaluating changes in the seasonal cycle of precipitation have led to the attribution of shifts in seasonal rainfall patterns in Southern Madagascar to climate change Rigden et al. (2024).
@@
An important issue raised by this work is the discrepancy between the modeled and observed findings, as seen, for example, in the deviance statistic results for the observations versus CESM Large Ensemble (Figs. 2b, 6b, 8b). This may be due to a particular sequence of internal variability in the observations, measurement error in the observational record, or issues with the simulation of extreme weather in the CESM Large Ensemble. It would be helpful to test if these discrepancies also occur with other models. The use of multiple models by the WWA (Philip et al. (2020), OK2021), may help mitigate against this kind of problem.
@@
Overall, our results suggest uncertainty in the interpretation of the dependence of the parameters of extreme event distributions on the global mean surface temperature as reflecting ACC. The proposed and demonstrated deviance statistic, MRL plot, and a careful analysis of uncertainty ranges, especially for $\alpha$, may help avoid misguided confidence in attribution results. As noted in the introduction, WWA does not exclusively rely upon the empirical methodology evaluated here in order to reach conclusions, and also evaluates other historical data, model simulations, and the dynamical context. Philip et al. (2020) also highlight some caveats associated with their study, including the necessity of testing for other probability laws (e.g., Gaussian, Gamma, etc.) and covariates (ENSO, PDO, etc.), among other issues. Nevertheless, the observations-based component of the WWA methodology examined here still figures prominently in their analyses, and we have highlighted several caveats with its interpretation, foremost that internal variability can give rise to the appearance of strong covariance between GMST and local extremes. It is important to study when the ACC signal emerges distinctly relative to internal variability on a case-by-case basis. The ACC signal is known to have emerged in most regions of the world for temperature extremes, while this may not be the case for variables like precipitation that have much higher variability. Time of emergence is a factor that should always be considered when discussing the results of an attribution study. Our results do not provide evidence against a relationship between GMST and local extremes but, rather, that internal variability needs to be taken into account when fitting a distribution function to the record. As currently implemented, the WWA approach can lead to first-order biases in some cases. It, therefore, seems that it may be best to use a multi-method approach to ensure a robust attribution approach, including the storyline approach that can add insight about the change to extreme events (e.g., Shepherd et al., 2018; Leach et al., 2024; Robin and Ribes, 2020b; Sippel et al., 2024).
@@
We suggest two key takeaways. First, controlling for internal variability, possibly by making the distribution parameters depend on additional covariates such as ENSO, appears important in order to accurately estimate the mean ACC trend. This improvement upon the current method should be further developed. Inclusion of additional covariates is vital as a way to avoid biases in the estimate of ACC effects on extreme event statistics. Second, empirical attribution techniques should be examined by applying the approach to simulations with and without changes to anthropogenic emissions. Application to such simulations would make it much easier to identify the likely influence of other factors on the local response in the absence of ACC, and so to ensure that all relevant modes are included in the statistical model, assuming the models are right in this respect.